# MGMCpaper.bib

### v2

@@ 4.4.1 Approximation property for the perturbed problem
Assumption 2.2 assumes a regularity of the solution of the unperturbed problem (2.17). It turns out that the solution of the perturbed problem (4.36) has the same regularity.
@@
If $\mathcal{A}$ satisfies Assumption 2.2, then so does the perturbed operator $\widetilde{\mathcal{A}}=\mathcal{A}+\mathcal{B}^{*}\Gamma^{-1}\mathcal{B}$ and
@@
where $\Psi(\ell)$ is defined in Assumption 2.2.
@@
The solution $u\in V$ of (4.36) satisfies $a(u,\varphi)=\langle\tilde{f},\varphi\rangle_{H}$ for any $\varphi\in V$, where we let $\tilde{f}:=f-\mathcal{B}^{*}\Gamma^{-1}\mathcal{B}u$. Hence, Assumption 2.2 for $\mathcal{A}$ together with $\|u\|_{H}\leq C\|u\|_{V}\leq C^{\prime}\|f\|_{H}$ implies
@@
The bound (4.39) is obtained by adapting Proposition 2.2 to $\widetilde{\mathcal{A}}$, again under Assumption 2.2. ∎
@@
Assumption 4.1.
There exists a constant $C_{\Psi}>1$, such that the function $\Psi\colon\mathbb{N}_{0}\to[0,\infty)$ in Assumption 2.2 satisfies
@@
$\|\varphi_{\ell}\|_{V}\leq\|\varphi_{\ell}\|_{H}/\sqrt{\Psi(\ell)},\quad\text{for all $\varphi_{\ell}\in V_{\ell}\,$}.$ (4.41)
@@
Let Assumptions 2.1 and 4.1 hold. Then, there exists a constant $C>0$ such that
@@
Let $x,y\in\mathbb{R}^{n_{\ell}}$. Then it follows from Assumptions 2.1 and 4.1 that
@@
For example, piecewise polynomial FE spaces $V_{\ell}$ on quasi-uniform triangulations, with $V=H^{1}(D)$ and $H=L^{2}(D)$, satisfy the inverse estimate with $\sqrt{\Psi(\ell)}\asymp h_{\ell}$; see e.g., [Quarteroni.A_Vali_book_1994_NAPDE, Prop. 6.3.2].
@@
Let Assumptions 2.1 and 4.1 be satisfied. Then, there exists a constant $\tilde{C}>0$ such that
@@
Let Assumptions 2.1, 2.2 and 4.1 hold. Then, we have the approximation property for the perturbed matrix $\widetilde{A}_{\ell}$
@@
Let Assumption 2.1 hold. Then, $A_{\ell}$ as in (2.9) satisfies
@@
[…]
@@
Let Assumption 2.1 hold. Suppose $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for all $\ell=0,\dots,L$, with $C_{A}>0$ independent of $\ell$. Then, the symmetric Gauss–Seidel iteration matrix $\widetilde{M}_{\ell}^{\text{(SGS)}}$ as in (4.34) satisfies
@@
Noting that Assumption 2.1 implies $\|B_{\ell}\Gamma^{-1}B_{\ell}^{\top}\|_{2}\leq\|\Gamma^{-1}\|_{2}\|\mathcal{B}\|_{H\to\mathbb{R}^{\beta}}^{2}(\Phi(\ell))^{2}$, we use the assumption $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$, and Lemma 4.18 to obtain
@@
In the lemma above we assumed $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for the unpurterbed precision matrix; this is a standard assumption (see e.g. [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.30]).
@@
Let Assumptions 2.1, 2.2 and 4.1 hold and suppose $y\in\mathbb{R}^{\beta}$. Consider the MGMC update for $\mathcal{N}(\widetilde{\mu}_{L},\widetilde{A}_{L}^{-1})$ with mean and covariance given by $\widetilde{\mu}_{L}=A_{L}^{-1}B_{L}(\Gamma+B_{L}^{\top}A_{L}^{-1}B_{L})^{-1}y$ and $\widetilde{A}_{L}^{-1}=(A_{L}+B_{L}\Gamma^{-1}B_{L}^{\top})^{-1}$. Assume further that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with moments $\mathbb{E}[\theta_{L}^{(0)}]$ and $\mathrm{Cov}(\theta_{L}^{(0)})$ that are bounded such that
$\max\{||\widetilde{A}_{L}^{1/2}\mathbb{E}[\theta_{L}^{(0)}]||_{2},||\widetilde{A}_{L}^{1/2}\mathrm{Cov}(\theta_{L}^{(0)})\widetilde{A}_{L}^{1/2}||_{2}\}<C_{0}$
@@
From the discussions in this section, Proposition 4.5 holds for the MGMC for $\mathcal{N}(\widetilde{\mu}_{L},\widetilde{A}_{L}^{-1})$ and thus $\|\tilde{X}_{L}\|_{\tilde{A}_{L}}\leq q<1$ holds with an $L$-independent constant $q$. To prove the statement we follow the argument in the proofs of Theorems 4.12 and 4.13. Since the moments of the initial condition is assumed to be bounded, it suffices to bound $\|F_{L}^{\top}\widetilde{A}_{L}^{-1/2}\|_{2}$ and $\|\widetilde{A}_{L}^{-1/2}f_{L}\|_{2}$.
@@
with $v:=(\Gamma+B_{L}^{\top}A_{L}^{-1}B_{L})^{-1}y_{L}$. Assumption 2.1 on $P_{L}$ implies
@@
where $C>0$ depends on the norm $\|\mathcal{B}\|_{H\to\mathbb{R}^{\beta}}$, the constants $c_{1}$ and $c_{2}$ in Assumption 2.1, and the smallest eigenvalues of $\mathcal{A}$ and $\Gamma^{-1}$ but independent of $L$. ∎
@@ 4.5 Optimality of MGMC
Under Assumption 2.2, the sequence $(F_{L}^{\top}\theta_{L}^{*})_{L\in\mathbb{N}}$ with $\theta_{L}^{*}\sim\mathcal{N}(\widetilde{A}_{L}^{-1}f_{L},\widetilde{A}_{L}^{-1})$ converges in distribution to the $\mathbb{R}$-valued Gaussian random variable $(\widetilde{v},\mathcal{F})$ with mean and covariance given in Eqs. (2.21)–(2.22) (but with $\chi=\psi=\mathcal{F}$).
@@
Perhaps not very surprisingly, the notion of optimality of MGMC in Corollary 4.21 is closely related to that of multigrid solvers. For this, consider for example [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.16] and [Hackbusch.W_2016_IterativeSolutionLarge, Remark 11.17] which shows that under certain conditions on the cycle parameter $\gamma$ the cost per multigrid cycle is bounded by a constant times the problem size $n_{L}$. Further, in [Hackbusch.W_2016_IterativeSolutionLarge, Section 11.5.5] the application of a nested multigrid iteration to PDE based problems is considered. It is shown there that this approach can reduce the error to the order of the discretisation error at a cost that can be bounded by a constant times $n_{L}$.
@@ 5.1.1 Prior distribution
As the prior, we consider $d$-dimensional random Gaussian processes on the unit cube $\Omega=[0,1]^{d}$ with mean zero and the following two covariance operators as concrete examples of the operator $\mathcal{A}$ in Section 2.1:
@@
with the homogeneous Dirichlet boundary condition $u(x)=0$ for $x\in\partial\Omega$ and
[…]
with $u(x)=\partial{u}/\partial n(x)=0$ for $x\in\partial\Omega$. In the numerical experiments we fix the correlation length $\kappa^{-1}=0.1$ in two dimensions and set $\kappa^{-1}=1.0$ in three dimensions unless stated otherwise. Observe that if the operators were defined in the domain $\Omega=\mathbb{R}^{d}$ instead of the unit cube, then the covariance function of the shifted Laplace operator $\mathcal{A}^{\mathrm{(SL)}}$ in (5.2) would be of Matérn class $\nu=-\frac{1}{2}$ in $d=3$ dimensions and $\nu=0$ in $d=2$ dimensions; for the squared shifted Laplace operator $\mathcal{A}^{(\mathrm{SSL})}$ in (5.3) the Matérn class would be $\nu=1$. For short distances $||x-y||_{2}\ll\kappa^{-1}$ the covariance function of the problem on $\mathbb{R}^{d}$ satisfies
$\mathrm{Cov}[\phi(x),\phi(y)]\propto\begin{cases}z^{\nu}K_{\nu}(\kappa r)\\ K_{0}(z)\\ z^{-\frac{1}{2}}K_{\frac{1}{2}}(z)\end{cases}=\begin{cases}\mathcal{O}(1)&\text{for $\nu>0$}\\ \mathcal{O}(-\log(z))&\text{for $\nu=0$}\\ \mathcal{O}(z^{-1})&\text{for $\nu=-\frac{1}{2}$}\end{cases}\qquad\text{with $z:=\kappa||x-y||_{2}\ll 1$.}$ (5.4)
@@
Since it is not possible to sample from an infinite dimensional prior, a grid is constructed by dividing the domain into $n_{L}^{d}$ quadrilateral cells of size $h_{L}^{d}$ and the second order operator in (5.2) is discretised with both a simple finite difference discretisation (FD) and with a lowest order conforming piecewise (multi-)linear finite element (FEM) discretisation on this grid in $d=2,3$ dimensions. For the fourth order operator $\mathcal{A}^{\mathrm{(SSL)}}$ in (5.3) we only consider the case $d=2$ and use the 13-point FD discretisation that is written down for example in [bjorstad1983fast]. The homogeneous Dirichlet boundary condition $u(x)=0$ is enforced by implicitly setting the solution on the boundary to zero and only storing the $(n_{L}-1)^{d}$ unknowns associated with the interior vertices. For the operator $\mathcal{A}^{\mathrm{(SSL)}}$ we enforce the second boundary condition $\partial u/\partial n(x)=0$ as described in [bjorstad1983fast].
@@ 5.1.3 Samplers
To generate a new state $\theta_{L}^{(m+1)}$ from the current state $\theta_{L}^{(m)}$ we use the multigrid Monte Carlo update in Alg. 1 with $\nu_{1}=1$ forward Gibbs-sweeps and $\nu_{2}=1$ backward Gibbs sweeps on each level, where each sweep consists of an iteration over the entire lattice. When sampling from the posterior we always use the Gibbs-sampler with low-rank updates as written down in Alg. 3. The number of levels is chosen such that the coarsest lattice consists of $n_{0}^{d}$ cells where $n_{0}$ is either an odd number (for example $n_{0}=3$ if $n_{L}=48=2^{4}\cdot 3$) or $n_{0}=2$ and hence there is only a very small number of interior grid points on this level. On the very coarsest level a few iterations of the symmetric Gibbs sampler with low-rank update are applied. Multilinear interpolation is used between different levels of the multigrid hierarchy. In a finite element setting this implies that the function spaces are nested. For the prior covariance in (5.2) we employ a V-cycle but a W-cycle is used for the prior defined by (5.3).
@@
for $\theta_{L}^{(m)}$ where $g_{L}$ is the solution of the triangular system $U_{L}^{\top}g_{L}=\mathfrak{P}_{L}f_{L}$. If $\widetilde{A}_{L}$ is sparse, the permutation $\mathfrak{P}_{L}$ can be chosen such as to minimise the number of non-zero entries in $U_{L}$ which is crucial to make the triangular solve in (5.7) efficient. In our implemenation we use the Simplicial Cholesky factorisation from the widely used Eigen library [eigenweb]. We find that for the problems considered here this gives slightly better performance than the implementation in the CholMod package [chen2008algorithm].
@@ 5.2 Performance
We start by empirically confirming the cost analysis in Section 3.6 and investigating the grid-independence of the IACT stated in Theorem 4.20. For this, we measure the time for producing a single Monte Carlo update (with the Gibbs- and Multigrid MC algorithm) and for drawing an independent sample with the Cholesky sampler. Tab. 1 shows these results in $d=2$ and $d=3$ dimensions for different priors, we consider both the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in (5.2) and its square $\mathcal{A}^{(\mathrm{SSL})}=(-\Delta+\kappa^{2}I)^{2}$ in (5.3). For the Gibbs and MGMC samplers we also list the IACT (which is 1 for Cholesky). As expected from (3.16) and Theorem 3.2, the time per sample grows approximately in proportion to the problem size for Multigrid MC and the Gibbs sampler. For the Cholesky sampler the growth in runtime is more rapid, in particular in $d=3$ dimensions: going from the $48\times 48\times 48$ to the $64\times 64\times 64$ lattice, the number of unknowns increases by a factor $(63/47)^{3}\approx 2.4$ but the cost of the Cholesky sampler is 3.8×3.8$\times$3.8 × larger. To account for the fact that the samples in the Markov chain are correlated while the Cholesky sampler produces independent samples, we multiply the time per sample with the IACT for the Gibbs- and Multigrid MC samplers. The resulting time per independent sample is shown in the final two columns of Tab. 1, and we use this number for a fair comparison with the Cholesky sampler. One application of the Gibbs sampler is slightly cheaper than a Multigrid MC update for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$, whereas for $\mathcal{A}^{(\mathrm{SSL})}$ the difference if more pronounced and MGMC is more than twice as expensive as Gibbs. As predicted by Theorem 4.20, the IACT for MGMC is roughly independent of the resolution. For the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$ it lies 1.1 and 1.4, whereas it is between 2.2 and 4.0 for $\mathcal{A}^{(\mathrm{SSL})}$. This is in stark contrast to the Gibbs sampler, where the IACT grows rapidly and can in fact not be reliably estimated on the finer lattices. Despite being cheaper if a single update is considered, the large IACT means that the Gibbs sampler is not competitive overall: on the finest lattices we considered, producing an independent sample is several orders of magnitude more expensive than with the other two methods. Comparing Cholesky and MultigridMC, the latter is roughly a factor two slower in $d=2$ dimensions for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$, for $\mathcal{A}^{(\mathrm{SSL})}$ the difference is even more pronounced with Cholesky being about five times as fast. However, MGMC is significantly faster for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$ in $d=3$ dimensions, in particular for finer resolutions. This can be attributed to the fact that – in contrast to Cholesky – one Multigrid MC update incurs a cost that can be bounded linearly in the number of unknowns and to the grid-independent IACT of Multigrid MC (see also Fig. 3, right).

### v3

@@ 4.4.1 Approximation property for the perturbed problem
Assumption 2 assumes a regularity of the solution of the unperturbed problem (2.17). It turns out that the solution of the perturbed problem (4.36) has the same regularity.
@@
If $\mathcal{A}$ satisfies Assumption 2, then so does the perturbed operator $\widetilde{\mathcal{A}}=\mathcal{A}+\mathcal{B}^{*}\Gamma^{-1}\mathcal{B}$ and
@@
where $\Psi(\ell)$ is defined in Assumption 2.
@@
The solution $u\in V$ of (4.36) satisfies $a(u,\varphi)=\langle\tilde{f},\varphi\rangle_{H}$ for any $\varphi\in V$, where we let $\tilde{f}:=f-\mathcal{B}^{*}\Gamma^{-1}\mathcal{B}u$. Hence, Assumption 2 for $\mathcal{A}$ together with $\|u\|_{H}\leq C\|u\|_{V}\leq C^{\prime}\|f\|_{H}$ implies
@@
The bound (4.39) is obtained by adapting Proposition 2.2 to $\widetilde{\mathcal{A}}$, again under Assumption 2. ∎
@@
Assumption 3.
There exists a constant $C_{\Psi}>1$, such that the function $\Psi\colon\mathbb{N}_{0}\to[0,\infty)$ in Assumption 2 satisfies
@@
[…]
@@
Let Assumptions 1 and 3 hold. Then, there exists a constant $C>0$ such that
@@
Let $x,y\in\mathbb{R}^{n_{\ell}}$. Then it follows from Assumptions 1 and 3 that
@@
For example, piecewise polynomial FE spaces $V_{\ell}$ on quasi-uniform triangulations, with $V=H^{1}(D)$ and $H=L^{2}(D)$, satisfy the inverse estimate with $\sqrt{\Psi(\ell)}\asymp h_{\ell}$; see e.g., [undefaao, Prop. 6.3.2].
@@
Let Assumptions 1 and 3 be satisfied. Then, there exists a constant $\tilde{C}>0$ such that
@@
Let Assumptions 1, 2 and 3 hold. Then, we have the approximation property for the perturbed matrix $\widetilde{A}_{\ell}$
@@
Let Assumption 1 hold. Then, $A_{\ell}$ as in (2.9) satisfies
@@
Let $x\in\mathbb{R}^{n_{\ell}}$. By the definition (2.9) of $A_{\ell}$ we have $x^{\top}A_{\ell}x=\langle\mathcal{A}^{1/2}P_{\ell}x,\mathcal{A}^{1/2}P_{\ell}x\rangle_{H}=\|P_{\ell}x\|_{V}^{2}$. Thus, Assumption 1 implies
@@
Let Assumption 1 hold. Suppose $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for all $\ell=0,\dots,L$, with $C_{A}>0$ independent of $\ell$. Then, the symmetric Gauss–Seidel iteration matrix $\widetilde{M}_{\ell}^{\text{(SGS)}}$ as in (4.34) satisfies
@@
Noting that Assumption 1 implies $\|B_{\ell}\Gamma^{-1}B_{\ell}^{\top}\|_{2}\leq\|\Gamma^{-1}\|_{2}\|\mathcal{B}\|_{H\to\mathbb{R}^{\beta}}^{2}(\Phi(\ell))^{2}$, we use the assumption $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$, and Lemma 4.20 to obtain
@@
In the lemma above we assumed $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for the unpurterbed precision matrix; this is a standard assumption (see e.g. [undefah, Theorem 11.30]).
@@
Let Assumptions 1, 2 and 3 hold and suppose $y\in\mathbb{R}^{\beta}$. Consider the MGMC update for $\mathcal{N}(\widetilde{\mu}_{L},\widetilde{A}_{L}^{-1})$ with mean and covariance given by $\widetilde{\mu}_{L}=A_{L}^{-1}B_{L}(\Gamma+B_{L}^{\top}A_{L}^{-1}B_{L})^{-1}y$ and $\widetilde{A}_{L}^{-1}=(A_{L}+B_{L}\Gamma^{-1}B_{L}^{\top})^{-1}$. Assume further that the initial sample $\theta_{L}^{(0)}$ is drawn from a distribution with moments $\mathbb{E}[\theta_{L}^{(0)}]$ and $\mathrm{Cov}(\theta_{L}^{(0)})$ that are bounded such that
[…]
@@
From the discussions in this section, Proposition 4.7 holds for the MGMC for $\mathcal{N}(\widetilde{\mu}_{L},\widetilde{A}_{L}^{-1})$ and thus $\|\tilde{X}_{L}\|_{\tilde{A}_{L}}\leq q<1$ holds with an $L$-independent constant $q$. To prove the statement we follow the argument in the proofs of Theorems 4.14 and 4.15. Since the moments of the initial condition is assumed to be bounded, it suffices to bound $\|F_{L}^{\top}\widetilde{A}_{L}^{-1/2}\|_{2}$ and $\|\widetilde{A}_{L}^{-1/2}f_{L}\|_{2}$.
@@
with $v:=(\Gamma+B_{L}^{\top}A_{L}^{-1}B_{L})^{-1}y_{L}$. Assumption 1 on $P_{L}$ implies
@@
where $C>0$ depends on the norm $\|\mathcal{B}\|_{H\to\mathbb{R}^{\beta}}$, the constants $c_{1}$ and $c_{2}$ in Assumption 1, and the smallest eigenvalues of $\mathcal{A}$ and $\Gamma^{-1}$ but independent of $L$. ∎
@@ 4.5 Optimality of MGMC
[…]
@@
Perhaps not very surprisingly, the notion of optimality of MGMC in Corollary 4.23 is closely related to that of multigrid solvers. For this, consider for example [undefah, Theorem 11.16] and [undefah, Remark 11.17] which shows that under certain conditions on the cycle parameter $\gamma$ the cost per multigrid cycle is bounded by a constant times the problem size $n_{L}$. Further, in [undefah, Section 11.5.5] the application of a nested multigrid iteration to PDE based problems is considered. It is shown there that this approach can reduce the error to the order of the discretisation error at a cost that can be bounded by a constant times $n_{L}$.
@@ 5.1.1 Prior distribution
As the prior, we consider $d$-dimensional random Gaussian processes on the unit cube $D=[0,1]^{d}$ with mean zero and the following two covariance operators as concrete examples of the operator $\mathcal{A}$ in Section 2.1:
$\mathcal{A}^{\mathrm{(SL)}}=-\Delta+\kappa^{2}I\qquad\text{(Shifted Laplace)}$ (5.2)
with the homogeneous Dirichlet boundary condition $u(x)=0$ for $x\in\partial D$ and
$\mathcal{A}^{\mathrm{(SSL)}}=(-\Delta+\kappa^{2}I)^{2}=\Delta^{2}-2\kappa^{2}\Delta+\kappa^{4}I\qquad\text{(Squared Shifted Laplace)}$ (5.3)
with $u(x)=\partial{u}/\partial n(x)=0$ for $x\in\partial D$. In the numerical experiments we fix the correlation length $\kappa^{-1}=0.1$ in two dimensions and set $\kappa^{-1}=1.0$ in three dimensions unless stated otherwise. Observe that if the operators were defined in the domain $D=\mathbb{R}^{d}$ instead of the unit cube, then the covariance function of the shifted Laplace operator $\mathcal{A}^{\mathrm{(SL)}}$ in (5.2) would be of Matérn class $\nu=-\frac{1}{2}$ in $d=3$ dimensions and $\nu=0$ in $d=2$ dimensions; for the squared shifted Laplace operator $\mathcal{A}^{(\mathrm{SSL})}$ in (5.3) the Matérn class would be $\nu=1$. For short distances $x-y||_{2}\ll\kappa^{-1}$ the covariance function of the problem on $\mathbb{R}^{d}$ satisfies
@@
Since sampling from an infinite dimensional prior is impossible, by dividing the domain into $n_{L}^{d}$ quadrilateral cells of size $h_{L}^{d}$ a grid is constructed and the second order operator in (5.2) is discretised with both a simple finite difference discretisation (FD) and with a lowest order conforming piecewise (multi-)linear FE discretisation on this grid in $d=2,3$ dimensions. While the theory in Section 4 was derived in the FE framework, we include the FD discretisation since this is often used in practice and to demonstrate that our results also apply in this setting. For the fourth order operator $\mathcal{A}^{\mathrm{(SSL)}}$ in (5.3) we only consider the case $d=2$ and use the 13-point FD discretisation that is written down for example in [undefaat]. Note that the FE discretisation of the biharmonic operator is significantly more involved and requires bespoke multigrid patch-smoothers (see e.g. [undefaau]), which is why we do not consider it here. The homogeneous Dirichlet boundary condition $u(x)=0$ is enforced by implicitly setting the solution on the boundary to zero and only storing the $(n_{L}-1)^{d}$ unknowns associated with the interior vertices. For the operator $\mathcal{A}^{\mathrm{(SSL)}}$ we enforce the second boundary condition $\partial u/\partial n(x)=0$ as described in [undefaat].
The convergence theory in Section 4 can in fact be easily adapted to include general boundary conditions, provided the corresponding multigrid solver converges for these boundary conditions.
@@ 5.1.3 Samplers
To generate a new state $\theta_{L}^{(m+1)}$ from the current state $\theta_{L}^{(m)}$ we use the multigrid Monte Carlo update in Alg. 1 with $\nu_{1}=1$ forward Gibbs-sweeps and $\nu_{2}=1$ backward Gibbs sweeps on each level, where each sweep consists of an iteration over the entire lattice. When sampling from the posterior we always use the Gibbs-sampler with low-rank updates as written down in Alg. 3. The number of levels is chosen such that the coarsest lattice consists of $n_{0}^{d}$ cells where $n_{0}$ is either an odd number (for example $n_{0}=3$ if $n_{L}=48=2^{4}\cdot 3$) or $n_{0}=2$ and hence there is only a very small number of interior grid points on this level. On the very coarsest level a few iterations of the symmetric Gibbs sampler with low-rank update are applied. In FE methods, the nesting of the function spaces $V_{\ell-1}\subset V_{\ell}$ induces the prolongation, which is equivalent to a multilinear interpolation between the nodal degrees of freedom. We use the same multilinear intergrid operator for the finite difference discretisation. For the prior covariance in (5.2) we employ a V-cycle but a W-cycle is used for the prior defined by (5.3).
@@
for $\theta_{L}^{(m)}$ where $g_{L}$ is the solution of the triangular system $U_{L}^{\top}g_{L}=\mathfrak{P}_{L}f_{L}$. If $\widetilde{A}_{L}$ is sparse, the permutation $\mathfrak{P}_{L}$ can be chosen such as to minimise the number of non-zero entries in $U_{L}$ which is crucial to make the triangular solve in (5.7) efficient. In our implemenation we use the Simplicial Cholesky factorisation from the widely used Eigen library [undefaav]. We find that for the problems considered here this gives slightly better performance than the implementation in the CholMod package [undefaaw].
@@ 5.2 Performance
We start by empirically confirming the cost analysis in Section 3.6 and investigating the grid-independence of the IACT stated in Theorem 4.22. For this, we measure the time for producing a single Monte Carlo update (with the Gibbs- and Multigrid MC algorithm) and for drawing an independent sample with the Cholesky sampler. Tab. 1 shows these results in $d=2$ and $d=3$ dimensions for different priors, we consider both the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in (5.2) and its square $\mathcal{A}^{(\mathrm{SSL})}=(-\Delta+\kappa^{2}I)^{2}$ in (5.3). For the Gibbs and MGMC samplers we also list the IACT (which is 1 for Cholesky). As expected from (3.17) and Theorem 3.2, the time per sample grows approximately in proportion to the problem size for Multigrid MC and the Gibbs sampler. For the Cholesky sampler the growth in runtime is more rapid, in particular in $d=3$ dimensions: going from the $48\times 48\times 48$ to the $64\times 64\times 64$ lattice, the number of unknowns increases by a factor $(63/47)^{3}\approx 2.4$ but the cost of the Cholesky sampler is $3.8\times$ larger. To account for the fact that the samples in the Markov chain are correlated while the Cholesky sampler produces independent samples, we multiply the time per sample with the IACT for the Gibbs- and Multigrid MC samplers. The resulting time per independent sample is shown in the final two columns of Tab. 1, and we use this number for a fair comparison with the Cholesky sampler. One application of the Gibbs sampler is slightly cheaper than a Multigrid MC update for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$, whereas for $\mathcal{A}^{(\mathrm{SSL})}$ the difference if more pronounced and MGMC is more than twice as expensive as Gibbs. As predicted by Theorem 4.22, the IACT for MGMC is roughly independent of the resolution. For the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$ it lies 1.1 and 1.4, whereas it is between 2.2 and 4.0 for $\mathcal{A}^{(\mathrm{SSL})}$. This is in stark contrast to the Gibbs sampler, where the IACT grows rapidly and can in fact not be reliably estimated on the finer lattices. Despite being cheaper if a single update is considered, the large IACT means that the Gibbs sampler is not competitive overall: on the finest lattices we considered, producing an independent sample is several orders of magnitude more expensive than with the other two methods. Comparing Cholesky and MGMC, the latter is roughly a factor two slower in $d=2$ dimensions for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$, for $\mathcal{A}^{(\mathrm{SSL})}$ the difference is even more pronounced with Cholesky being about five times as fast. However, MGMC is significantly faster for the shifted Laplace operator $\mathcal{A}^{(\mathrm{SL})}$ in $d=3$ dimensions, in particular for finer resolutions. This can be attributed to the fact that – in contrast to Cholesky – one Multigrid MC update incurs a cost that can be bounded linearly in the number of unknowns and to the grid-independent IACT of Multigrid MC (see also Fig. 3, right).
shifted Laplace $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in $d=2$ dimensions, FEM discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$32\times 32$ $3.1\pm 0.4$ $1.12\pm 0.12$ $0.15$ $0.22$ $0.08$ $0.46$ $0.25$
$64\times 64$ $0.5\pm 2.1 1.1$$13\pm 0.12$ $0.62$ $0.91$ $0.39$ $6.47$ $1.04$
$128\times 128$ $47.3\pm 16.4$ $1.15\pm 0.13$ $2.53$ $3.71$ $1.79$ $119.93$ $4.28$
$256\times 256$ $95.1\pm 42.2$ $1.18\pm 0.14$ $10.18$ $15.17$ $9.14$ $968.49$ $17.96$
$512\times 512$ $264.2\pm 169.1$ $1.21\pm 0.15$ $43.03$ $66.69$ $47.23$ $11370.84$ $80.53$
shifted Laplace $\mathcal{A}^{(\mathrm{SL})}=-\Delta+\kappa^{2}I$ in $d=3$ dimensions, FD discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$16\times 16\times 16$ $2.6\pm 0.3$ $1.32\pm 0.19$ $0.57$ $0.74$ $0.48$ $1.47$ $0.98$
$32\times 32\times 32$ $4.7\pm 0.7$ $1.20\pm 0.14$ $5.20$ $7.62$ $12.58$ $24.60$ $9.17$
$48\times 48\times 48$ $10.3\pm 2.0$ $1.26\pm 0.17$ $24.87$ $30.97$ $82.81$ $257.20$ $39.17$
$64\times 64\times 64$ $20.6\pm 5.3$ $1.28\pm 0.17$ $50.11$ $67.89$ $317.23$ $1030.82$ $87.00$
squared shifted Laplace $\mathcal{A}^{(\mathrm{SSL})}=(-\Delta+\kappa^{2}I)^{2}$ in $d=2$ dimensions, FD discretisation
grid size IACT time / sample time / indep. sample
Gibbs MGMC Gibbs MGMC Chol. Gibbs MGMC
$32\times 32$ $2.4\pm 5.9 2.2$$22\pm 0.26$ $0.15$ $0.36$ $0.10$ $3.45$ $0.81$
$64\times 64$ $401.7\pm 4757.4 3.3$$35\pm 0.43$ $0.64$ $1.50$ $0.52$ $2167.88$ $5.02$
$128\times 128$ $1976.3\pm 2448.2$ $2.69\pm 0.35$ $2.61$ $6.15$ $2.49$ $5148.96$ $16.55$
$256\times 256$ $2573.9\pm 3415.3$ $3.23\pm 0.40$ $10.59$ $25.43$ $16.13$ $27265.49$ $82.01$
$512\times 512$ $1682.9\pm 1991.1$ $3.94\pm 0.57$ $45.09$ $110.39$ $81.68$ $75871.91$ $435.17$